COMPUTER NETWORKS

  1. What is the channel capacity if you have 16 signal levels and the bandwidth is 8 KHz?

  1. What is the wavelength if the speed/velocity of the wave is 100 m/s and frequency is 200Hz ?

  1. What is the frequency of the period is 4 s ?
Answer #1

Question 1

Channel capacity is calculated in bit rates. The formula for a noiseless channel is

Capacity = 2 * Bandwidth * log2 L

Kilo = 1000

L means the number of signal levels

Capacity = 2 * 8000 * log2 16 = 2 * 8000 * 4 = 64000 bps

Question 2

Speed = Wavelength x Wave Frequency

100 m/s = wavelength x 200 Hz

wavelength = 0.5 meters

Question 3

Frequency = 1/T Where T is period

Frequency = 1/4s = 0.25 Hz

1/s gives Hz unit and 1/4=0.25